A research group wants to determine whether the proportion of car accidents caused by drivers using cell phones has changed from the previous value of 13%. They obtained 10,000 auto accident reports and found that 17% were caused by drivers using cell phones. Computer the value of the test statistic.

options:

A. 16.533

B. 23.788

C. 4.044

D. 11.894
Answer #1

Solution :

H0 : p = 0.13

\hat p = 0.17

P0 = 0.13

1 - P0 = 1 - 0.13 = 0.87

Test statistic = z =

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.17 - 0.13/ [\sqrt(0.13 * 0.87) / 10000]

Test statistic = z = 11.894

correct option is = D